In a water solubility test of aqueous magnesium sulphate at room temperature, more than 25cm3 of baCl2 [aq] was added to its saturated solution. What is the maximum mass of aqueous MgSO4 dissolved in 100g of water at room temperature? [The density of saturated MgSO4 (aq) is 1.21gcm-3.]
1
Expert's answer
2020-12-21T04:17:04-0500

Solubility (MgSO4) = 35.1 g/(100 g water) (20 oC)

M(MgSO4) = 120 g/mol

M(MgSO4*7H2O) = 246 g/mol

Solubility (MgSO4*7H2O) = 35.1*246/120 = 72.0 g in 100-35.1/120*(246-120) g H2O

Solubility (MgSO4*7H2O) = 72 g in 63.1 g water

Solubility (MgSO4*7H2O) = 114g/(100g water)